OpenStudy (anonymous):

Subtract. (4x3 + 2x2 + 3x + 8) – (x3 – 5x2 + 11x – 9) 3x3 – 3x2 + 14x – 1 3x3 + 7x2 – 8x + 17 3x3 – 3x2 – 8x – 1 3x3 + 7x2 + 14x + 17

OpenStudy (zehanz):

You can only add/subtract like terms: 4x³-x³ = 3x³ 2x²-5x²=-3x² 3x-11x= ... 8-(-9)=... Try it. It is not that hard!